You are here: Foswiki>PartnerObserving Web>ItalyQueue2010>ObsInaf20100618 (22 Jun 2010, VincenzoTest)Edit Attach

INAF LBC+LUCIFER queue observing, 2010 June17-18

Observer: V. Testa, K.Boutsia
Telescope Operator: S. Allanson
Support Astronomer: O. Kuhn
Instrument: LUCIFER1 - LBC

Summary

Data Summary

Project Objects Filter ExpTime Notes
Mannucci HDFN 210zJHK H 2hr arc- flat- tell
Mannucci HDFN 210zJHK K 35min arc- flat- tell
Giallongo Q1623 Uspec, z-sloan 30min OB 1 redone in spec
Rigliaco Serpens Uspec, I-bessel 30 min  
Rigliaco Serpens B-bessel, V-bessel 10 min  
Rigliaco Serpens Uspec, R-bessel 20 min  
Rigliaco Serpens Uspec, B-bessel, V-bessel, R-bessel 5 min short frames

Details

02:30 - Opening the dome - starting interface

02:40 - Sky Flat fields

in the Ks band - imaging 0001-0036
  • 2secs / 8000-3000counts

in the J band 0037-0066
  • 2secs / 14000-3000counts

02:56 - TO unwraps azimuth so we won't hit limits

03:00 - Pointing Star near the field (WT10_385) 0067
  • Current IE = -84.5313
  • Current CA = -61.8785
  • Corrected IE = -50.0301
  • Corrected CA = -51.3113

03:03 - Persson standard BS9145 for collimation (POSITION,ACTIVE)

03:07 - Collimated! - preset to Telluric for Mannucci HDFN H
  • Guide star at the edge of field, moved out of position. Acquisition spolied!!

03:15 - Redo acquisition, star half out of the slit. Don't care, go on.
  • No way. Change guide star and repoint!!!!!!
  • Now it's ok,

03:22 - telluric Hnew 0073-0087

03:27 - Slew to HDFN for Mannucci H - acquisition-

03:41 - Mannucci_HDFN_H_3 0092-0103

04:47 - Y+1.5" offset, through slit to check
  • INITIALIZE

04:50 - Mannucci_HDFN_H_4 0105-0116

05:57 - Commit offset Y+1.5, change to K 0117-0118
  • through-slit, slight offset, give X+0.24
  • through-slit again with 10s * 2NDIT

06:03 - Mannucci_HDFN_K 0119-0127
  • below elev. 40 started vibrations in the Autoguider.
  • 8th integration already degraded 0126
  • stopped after 9th intergration
  • first 7 integrations are fine

06:55 - go to telluric, does not find.

06:56 - correct pointing on ACT0376 0129
  • Current IE = -50.0301
  • Current CA = -51.3113
  • Corrected IE = -64.4889
  • Corrected CA = -61.2641

07:03 - Slew to Telluric
  • had to change three guide stars, it's always a pain in the.....neck!
  • offset X+0.48

07:12 - Telluric! 0132-0155

07:19 - Arc & Flats
  • Arc H 0156-0159
  • Flat H 0160-0165
  • Arc K 0166-0169
  • Flat K 0170-0175

It took 45 minutes to do last telluric + calibration. This must be counted as science time to the PI

07:36 - SWITCH TO LBC

07:55 - Meanwhile 5Bias_Bino to warmup

08:00 - Connected Telescope
  • Slew to ACT0298 to correct pointing
  • NEW CA, IE: -79.7 -51.7
  • Corrected IE = -51.658227157489
  • Corrected CA = -79.716838767769

08:04 - Slew to Q1623_Uspec_1 (redo since it was observed out of specs)
  • dofpia
  • R -2902 293 -82 680 1147 216 -47 0.49 1.30
  • B -16143 0 0 0 0 0 0 0.00 6.36
  • R -304 -106 113 15 5 103 -15 0.46 0.16
  • B -4266 -381 8 1100 285 -91 -56 0.62 1.76
  • R 25 98 3 -36 -47 72 -12 0.47 0.10
  • B -688 -3 -357 147 21 101 -20 0.73 0.31
  • R -150 -157 141 74 33 83 -1 0.50 0.13
  • B -71 271 27 -21 -18 50 -13 0.83 0.12 a bit perfectionist, huh?
  • R -56 39 -254 -8 -3 24 -1 0.51 0.11
  • B 33 -267 -26 12 -65 -33 -21 0.77 0.12 stopped!

08:13 - Q1623_Uspec_1
  • 0.6" in Uspec, 0.5" in z-SLOAN!
  • Uspec is not able to do IA on line, focus degrades as OB goes by
  • z-SLOAN corrects the focus after every exposure and fwhm is wonderful throughout the OB

08:49 - Slew to Serpens
  • dofpia
  • R -673 9 -170 -139 -227 -30 6 0.54 0.31
  • B 80 -150 -511 -111 224 -449 -5 0.88 0.35
  • R 169 73 111 -16 36 56 -4 0.62 0.12
  • B 255 155 -244 -17 20 -84 -1 0.92 0.16
  • R -472 -160 39 -74 56 110 11 0.60 0.22
  • B 89 -63 164 79 -12 -13 -6 0.84 0.11

08:56 - Serpens_UsI_1 repeat since it was out of spec
  • Red out of focus immediately...
  • Not even found a Z4 correction to apply at the end of the exposure...
  • Stop and redo dofpia
  • R 851 -35 -214 324 -53 -344 0 0.55 0.44
  • B 1452 -122 -125 -24 -75 -471 8 0.84 0.65
  • R 58 132 253 -386 68 -33 -6 0.52 0.21
  • B 117 187 19 -92 15 -115 11 0.75 0.14
  • R -6 -70 -16 156 0 -21 9 0.45 0.11
  • B 351 -176 289 33 61 -126 -19 0.70 0.21

09:06 - Resend Serpens_UsI_1
  • Red still soon out of focus, sciencec frames are good in the center and blurred (comatic?) at the edges

09:36 - dofpia
  • R 436 79 -56 -631 -79 -196 -21 0.53 0.37
  • B 1668 232 -85 126 -141 -431 -16 0.77 0.72
  • R 48 32 261 290 67 56 9 0.65 0.18
  • B -5 -138 230 46 114 -59 9 0.95 0.13
  • R -201 86 -128 -127 62 2 8 0.69 0.14
  • B 122 35 -400 13 18 -117 -6 0.90 0.17

09:42 - Serpens_UsBVI_1_half (second half or yesterday night)
  • Still I-BESSEL with coma, perhaps filter offset to be redetermined?
  • We gave Z7=500 to the red to try to improve things, right main mirror failed to focus but goes on.
  • Z4+1000 to the red (09:58) it improved things a lot....
  • Z4+1000 @10:10

10:16 - Serpens_UsR_1_complete
  • to complete the OB interrupted on Jun 13 because of twilight

10:27 - dofpia
  • R -204 -83 68 -531 20 -126 -3 0.66 0.28
  • B 653 -220 238 -184 -48 -346 1 0.94 0.37
  • R 73 32 131 -195 101 0 16 0.69 0.14
  • B 43 64 -306 74 -23 -80 21 1.03 0.14
  • R 761 -177 -152 337 -282 -235 0 0.66 0.40
  • B 719 -60 72 72 -54 -271 9 1.00 0.34
  • R -484 91 106 -108 50 106 22 0.82 0.23
  • B -441 34 -202 -127 33 80 23 1.12 0.21
  • R 123 5 -91 -54 -22 -107 -10 0.65 0.12
  • B 488 261 190 93 -85 -168 -9 0.93 0.26 stopped

10:33 - Serpens_UsBVR_short
  • short exposures for calibration

10:41 - Slew to NGC6791
  • dofpia (6 iterations to converge, too late, sky is brightening move to standards)
  • serious problems in focusing tonight....

10:54 - SA110_UsBV_VRI

11:08 - SA113_UsBV_IVR

11:19 - Slew to FF position
  • Flats done: VV (10sec) - BR (2sec) -UsI (1sec) - UsZ (1sec)

11:55 - End of night - LUCIFER Darks

-- VincenzoTest - 17 Jun 2010
Topic revision: r5 - 22 Jun 2010, VincenzoTest
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback